# Counting rational points on Hirzebruch–Kleinschmidt varieties over global function fields

### v1

@@ 5.4. Decomposition of Hirzebruch–Kleinschmidt varieties
We now prove Theorem 1.2 from the introduction. We choose $L=-K_{X}=\gamma f+\xi f$ with $\gamma=r+1$ and $\xi=t-|\mathbf{a}|$. Hence, $A_{L}=B_{L}=1$, $a(L)=1$ and $b(L)=2$. Assume $a_{r}>0$. Then, the absolute convergence of $\zeta_{U,-K_{X}}(s)$ for $\Re(s)>1$ and the fact that $\zeta_{U,-K_{X}}(s)$ is a rational function in $q^{-s}$ follow directly from Theorem 5.2. Moreover, we have
$\lim_{s\to 1}(s-1)^{2}\zeta_{U,-K_{X}}(s)=C$
with $C$ given by (4). This proves the result in the case $a_{r}>0$. Now, assume $a_{r}=0$. On the one hand, Theorem 5.4 shows that $\zeta_{X,-K_{X}}(s)$ converges absolutely for $\Re(s)>1$, it is a rational function in $q^{-s}$ and
$\lim_{s\to 1}(s-1)^{2}\zeta_{X,-K_{X}}(s)=C$
with the same $C$ as before. On the other hand, by Lemma 5.15 we have
@@
$X\setminus U\simeq\left\{\begin{array}[]{ll}X_{d-1}(0,\ldots,0)\sqcup(\mathbb{P}^{r}_{1}\setminus\mathbb{P}^{r-1}_{1})\sqcup\bigg{(}\bigsqcup_{2\leq t^{\prime}<t}U_{t^{\prime}+r-1}(a_{1},\ldots,a_{r})\bigg{)}&\text{if }r>1,\\ \mathbb{P}^{t-1}_{2}\sqcup(\mathbb{P}^{1}_{1}\setminus\{P_{0}\})\sqcup\bigg{(}\bigsqcup_{2\leq t^{\prime}<t}U_{t^{\prime}+r-1}(a_{1})\bigg{)}&\text{if }r=1,\end{array}\right.$
with $X_{d-1}(0,\ldots,0)\simeq\mathbb{P}^{t-1}\times\mathbb{P}^{r-1}$ in the case $r>1$. Using Theorems 5.4 and 4.1, we see that
$\zeta_{X\setminus U,-K_{X}}(s)=\sum_{P\in X(K)\setminus U(K)}H(P)^{-s}$
converges absolutely for $\Re(s)>1$, is a rational function in $q^{-s}$ and has a simple pole at $s=1$. We conclude that
$\zeta_{U,-K_{X}}(s)=\zeta_{X,-K_{X}}(s)-\zeta_{X\setminus U,-K_{X}}(s)$
satisfies the desired properties. This completes the proof of Theorem 1.2.
6. Example: Hirzebruch surfaces
@@
In this last section we apply Theorems 4.1, 5.2 and 5.4 to the case of Hirzebruch surfaces $X:=X_{2}(a)$ with $a>0$ an integer, and general $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big. In this context, Lemma 5.15 gives
@@
with associated height zeta functions $\zeta_{\mathbb{P}^{1}}(\xi s),\zeta_{\mathbb{A}^{1}}(\gamma s)=\zeta_{\mathbb{P}^{1}}(\gamma s)-1$ and $\zeta_{U,L}(s)$ where $U:=U_{2}(a)$. Moreover, we have
$A_{L}=\frac{2}{\gamma},\quad B_{L}=\frac{a+2}{\gamma a+\xi},$
and the analytic properties of $\zeta_{U,L}(s)$ can be extracted from Theorem 5.2. In particular, $\zeta_{U,L}(s)$ converges absolutely for $\Re(s)>\max\{A_{L},B_{L}\}$, it has a pole at $s=\max\{A_{L},B_{L}\}$, and this pole is of order 2 if $A_{L}=B_{L}$ and of order 1 otherwise. For the zeta function $\zeta_{\mathbb{A}^{1}}(\gamma s)$, we know by Theorem 4.1 that it converges absolutely for $\Re(s)>A_{L}$ and it has a simple pole at $s=A_{L}$. Finally, the zeta function $\zeta_{\mathbb{P}^{1}}(\xi s)$ has no finite abscissa of absolute convergence if $\xi\leq 0$, and it converges absolutely for $\Re(s)>\frac{2}{\xi}$ with a simple pole at $s=\frac{2}{\xi}$ if $\xi>0$. This allows for a complete analysis of the analytic properties of the height zeta function of each component in (30).
In the case $L=-K_{X}$, we have $\gamma=2,\xi=2-a$ and $A_{L}=B_{L}=1$. As in Example 1.4, we find it necessary to remove the component $\mathbb{P}^{1}$ in (30) in order to verify the analogue of Conjecture 1.1 over global function fields. Indeed, the zeta series $\zeta_{X\setminus\mathbb{P}^{1}}(s)=\zeta_{\mathbb{A}^{1}}(\gamma s)+\zeta_{U,L}(s)$ converges absolutely for $\Re(s)>1$, it is a rational function in $q^{-s}$ and satisfies
$\lim_{s\to 1}(s-1)^{2}\zeta_{X\setminus\mathbb{P}^{1}}(s)=\lim_{s\to 1}(s-1)^{2}\zeta_{U,L}(s)=\frac{q^{4(1-g)}h_{K}^{2}}{\zeta_{K}(2)^{2}(a+2)2(q-1)^{2}\log(q)^{2}}.$
This is in accordance with [Bou02, Proposition 3.1] (see also p. 355 in loc. cit.) where the case of the anticanonical height zeta function over the good open subset $U_{2}(a)$ was studied.
Let us now choose $X=X_{2}(1)$ (i.e. $a=1$) for simplicity in order to illustrate further our results for more general big classes $L=\gamma h+\xi f\in\operatorname{Pic}(X)$. In the following table we denote by $\sigma_{1},\sigma_{2}$ and $\sigma_{3}$ the abscissas of absolute convergence of $\zeta_{\mathbb{P}^{1}}(\xi s),\zeta_{\mathbb{A}^{1}}(\gamma s)$ and $\zeta_{U,L}(s)$, respectively, for the first possible choices of $\gamma>0$ and $\xi>-\gamma$. We also record the order $b(L)$ of the pole of $\zeta_{U,L}(s)$ at $s=\sigma_{3}$.
$\gamma$ $\xi$ $\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $b(L)$
1 0 $\infty$ 2 3 1
1 1 2 2 2 1
1 2 1 2 2 1
2 $-1$infty 1 3 1
2 0 $\infty$ 1 $3/2$ 1
2 1 2 1 1 2
3 $-2$infty $2/3$ 3 1
3 $-1$infty $2/3$ $3/2$ 1
3 0 $\infty$ $2/3$ $1$ 1
The case $\gamma=\xi=1$ seems particularly interesting, since $\sigma_{1}=\sigma_{2}=\sigma_{3}=2$ and $b(L)=1$, meaning that the three components in (30) have height zeta functions with a simple pole at $s=2$. Moreover, we can compute
$\lim_{s\to 2}(s-2)\zeta_{\mathbb{P}^{1}}(s)=\lim_{s\to 2}(s-2)\zeta_{\mathbb{A}^{1}}(s)=\frac{h_{K}q^{2(1-g)}}{\zeta_{K}(2)(q-1)\log(q)},$
and using that $\xi>\frac{1}{2}\gamma$ we get
$\lim_{s\to 2}(s-2)\zeta_{U,L}(s)=\frac{h_{K}q^{2(1-g)}\mathbf{R}_{K}\left(-1,3\right)}{\zeta_{K}(3)\zeta_{K}(2)(q-1)\log(q)}.$
In the special case $K=\mathbb{F}_{q}(T)$ we can use that $g=0,h_{K}=1$ and
@@ 3 $-1$infty $2/3$ $3/2$ 1
(see e.g. [Ros02, p. 52]) and write
$\lim_{s\to 2}(s-2)\zeta_{\mathbb{P}^{1}}(s)=\lim_{s\to 2}(s-2)\zeta_{\mathbb{A}^{1}}(s)=\frac{q^{2}-1}{q\log(q)},$
and
$\lim_{s\to 2}(s-2)\zeta_{U,L}(s)=\frac{(q^{2}-1)(q^{2}+q+1)}{q^{2}\log(q)}.$
(see Remark 5.3(1)). Hence, in the case of the Hirzebruch surface $X_{2}(1)$ over $K=\mathbb{F}_{q}(T)$, with respect to the height function $H_{L}$ associated with $L=h+f\in\operatorname{Pic}(X_{2}(1))$, we can say that the component $U=U_{2}(1)$ in (30) has asymptotically more rational points of bounded height than the union $\mathbb{P}^{1}\sqcup\mathbb{A}^{1}$ of the other two components, since the residue of $\zeta_{U,L}(s)$ at $s=2$ is larger than the sum of the residues of $\zeta_{\mathbb{P}^{1}}(s)$ and $\zeta_{\mathbb{A}^{1}}(s)$.

### v2

@@ 5.5. Decomposition of Hirzebruch–Kleinschmidt varieties
We now prove Theorem 1.4 and Corollary 1.6 from the introduction. We choose $L=\eta_{X}^{-1}(-K_{X})=\gamma h+\xi f$ with $\gamma=\frac{r+1}{\eta_{X}}$ and $\xi=\frac{t-|\mathbf{a}|}{\eta_{X}}$ (see Proposition 3.2(2)). Hence, $a(L)=A_{L}=B_{L}=\eta_{X}$ and $b(L)=2$. Assume $a_{r}>0$. Then, the absolute convergence, rationality and holomorphicity statements about $\zeta_{U,-K_{X}}(s)=\zeta_{U,L}(\eta_{X}s)$ follow directly from Theorem 5.2. We also have
$\lim_{s\to 1}(s-1)^{2}\zeta_{U,-K_{X}}(s)=\eta_{X}^{-2}\lim_{s\to a(L)}(s-a(L))^{2}\zeta_{U,L}(s)$
and this value is exactly the one given in (4). Moreover, since $\zeta_{U,L}(s)$ has at most simple poles at each $s\in\mathscr{A}_{L}\cup\mathscr{B}_{L}$, we conclude that $\zeta_{U,-K_{X}}(s)$ has at most simple poles on $\Re(s)=1,s\not\in\left\{1+\frac{2\pi im}{\eta_{X}\log(q)}:m\in\mathbb{Z}\right\}$. This proves the result in the case $a_{r}>0$. Now, assume $a_{r}=0$. Then, by Lemmas 5.22 and 5.23 we have
@@
$\zeta_{U,-K_{X}}(s)=\left\{\begin{array}[]{ll}\zeta_{X,L}(\eta_{X}s)-\zeta_{X_{d-1}(a_{1},\ldots,a_{r-1}),L^{\prime}}(\eta_{X}s)&\text{if }r>1,\\ \zeta_{X,L}(\eta_{X}s)-\zeta_{\mathbb{P}^{t-1}}\left(ts\right)&\text{if }r=1,\end{array}\right.$
where in the case $r>1$ we put $L^{\prime}:=\gamma h^{\prime}+\xi f^{\prime}$ with $\{h^{\prime},f^{\prime}\}$ the basis of $\operatorname{Pic}(X_{d-1}(a_{1},\ldots,a_{r-1}))$ given in Proposition 3.2. In that case, we apply Theorem 5.4 to both $\zeta_{X,L}(s)$ and $\zeta_{X_{d-1}(a_{1},\ldots,a_{r-1}),L^{\prime}}(s)$ in order to deduce the desired properties of $\zeta_{U,-K_{X}}(s)$. Since $A_{L^{\prime}}=\frac{r}{r+1}\eta_{X}$ and $B_{L^{\prime}}=\eta_{X}$, we get $a(L^{\prime})=\eta_{X}$ and $b(L^{\prime})=1$. We deduce that both $\zeta_{X,L}(\eta_{X}s)$ and $\zeta_{X_{d-1}(a_{1},\ldots,a_{r-1}),L^{\prime}}(\eta_{X}s)$ converge absolutely for $\Re(s)>1$, are rational functions in $q^{-\eta_{X}s}$ and $\zeta_{X_{d-1}(a_{1},\ldots,a_{r-1}),L^{\prime}}(\eta_{X}s)$ has only simple poles for $\Re(s)=1$ while $\zeta_{X,L}(\eta_{X}s)$ has a double pole at $s=1$ and at most simple poles for $\Re(s)=1,s\not\in\left\{1+\frac{2\pi im}{\eta_{X}\log(q)}:m\in\mathbb{Z}\right\}$. Moreover, we see that $\zeta_{U,-K_{X}}(s)$ is holomorphic for
$\Re(s)>\frac{r}{r+1},\Re(s)\neq 1.$
Since $\frac{r}{r+1}\leq 1-\min\left\{\frac{1}{r+1},\frac{1}{t}\right\}$, we conclude that $\zeta_{U,-K_{X}}(s)$ satisfies all the desired analytic properties. Additionally, we have
$\lim_{s\to 1}(s-1)^{2}\zeta_{U,-K_{X}}(s)=\lim_{s\to 1}(s-1)^{2}\zeta_{X,-K_{X}}(s)=\eta_{X}^{-2}\lim_{s\to a(L)}(s-a(L))^{2}\zeta_{X,L}(s)$
and this values is the one given in (4). This proves Theorem 1.4 in the case $r\geq 1$. The case $r=1$ follows from Theorems 5.4 and 4.1 in an analogous way. We omit the details for brevity. This completes the proof of Theorem 1.4.
Finally, Corollary 1.6 is a direct consequence of Theorems 1.4 and 2.1.
6. Hirzebruch surfaces
@@
In this last section we apply our results to the case of Hirzebruch surfaces $X:=X_{2}(a)$ with $a>0$ an integer, and general $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big. In this context, Lemmas 5.20 and 5.23 give
@@
with associated height functions $H_{\mathbb{P}^{1}}^{\xi},H_{\mathbb{P}^{1}}^{\gamma}$ and $H_{L}$. Moreover, we have
$A_{L}=\frac{2}{\gamma},\quad B_{L}=\frac{a+2}{\gamma a+\xi}.$
Using the same notation as in Section 5.4, we get
$\tilde{N}(U,H_{L},q^{M})=Q_{L}(M)q^{a(L)M}+O\left(q^{\delta_{1}M}\right)\quad\text{as }M\to\infty,M\in\eta_{L}\mathbb{Z},$
for every $\delta_{1}>a^{\prime}(L)$, where $Q_{L}(M)=C_{1}M+C_{0}(M)$ with $C_{0}(M)\in\mathbb{R}$ depending only on $L$ and $M_{0}$ mod $\gamma_{0}(\gamma a+\xi)_{0}$ and
$C_{1}=\frac{q^{4(1-g)}\eta_{L}h_{K}^{2}}{\zeta_{K}(2)^{2}(\gamma a+\xi)\gamma(q-1)^{2}}$
if $A_{L}=B_{L}$, or $Q_{L}(M)>0$ depending only on $L$ and $M_{0}$ mod $(\gamma a+\xi)_{0}$ (resp. mod $\gamma_{0}$) if $A_{L}<B_{L}$ (resp. $A_{L}>B_{L}$).
Now, if $\xi\leq 0$ then the component $\mathbb{P}^{1}$ in (42) has no rational points of large height, and if $\xi>0$ then we have
$\tilde{N}(\mathbb{P}^{1},H_{\mathbb{P}^{1}}^{\xi},q^{M})=C_{2}q^{\frac{2M}{\xi}}+O\left(q^{\delta_{2}M}\right)\quad\text{as }M\to\infty,M\in\xi\mathbb{Z},$
for every $\delta_{2}>\frac{1}{2\xi}$, where
$C_{2}=\frac{h_{K}q^{2(1-g)}}{\zeta_{K}(2)(q-1)}.$
Finally, the component $\mathbb{A}^{1}$ in (42) satisfies
$\tilde{N}(\mathbb{A}^{1},H_{\mathbb{P}^{1}}^{\gamma},q^{M})=C_{2}q^{\frac{2M}{\gamma}}+O\left(q^{\delta_{3}M}\right)\quad\text{as }M\to\infty,M\in\gamma\mathbb{Z},$
for every $\delta_{3}>\frac{1}{2\gamma}$.
In the case $L=-K_{X}$, we have $\gamma=2,\xi=2-a,a(L)=1$ and $\eta_{L}=\mathrm{g.c.d.}(2,a)$. As in Example 1.7, we can distinguish two cases:
  1. (1)
    If $a=1$ then $\eta_{L}=1$ and
    $\tilde{N}(X,H_{-K_{X}},q^{M})\sim\tilde{N}(\mathbb{P}^{1},H_{\mathbb{P}^{1}},q^{M})\sim C_{2}q^{2M}\quad\text{as }M\to\infty,M\in\mathbb{Z}.$
  2. (2)
    If $a\geq 2$ then
    $\tilde{N}(X,H_{-K_{X}},q^{M})\sim\tilde{N}(U,H_{-K_{X}},q^{M})\sim C_{1}Mq^{M}\quad\text{as }M\to\infty,M\in\eta_{L}\mathbb{Z}.$
Finally, in order to show an example different from the anticanonical height, let us choose $a=\gamma=\xi=1$. Then, since $A_{L}=2,B_{L}=\frac{3}{2},a(L)=2,\eta_{L}=1$, we get
$\tilde{N}(U,H_{L},q^{M})\sim C_{3}q^{2M}\text{ and }\tilde{N}(\mathbb{P}^{1},H_{\mathbb{P}^{1}},q^{M})\sim\tilde{N}(\mathbb{A}^{1},H_{\mathbb{P}^{1}},q^{M})\sim C_{2}q^{2M}\quad\text{as }M\to\infty,M\in\mathbb{Z},$
where
$C_{3}=\frac{q^{2(1-g)}h_{K}}{\zeta_{K}(2)(q-1)}\sum_{D\geq 0}\tilde{N}_{1}(D)q^{-3\deg(D)}=\frac{q^{2(1-g)}h_{K}\mathbf{R}_{K}(-1,3)}{\zeta_{K}(2)\zeta_{K}(3)(q-1)}.$
In the special case $K=\mathbb{F}_{q}(T)$ we can use that $g=0,h_{K}=1,\mathbf{R}_{K}(-1,3)=q\zeta_{K}(2)$ and
@@
(see, e.g., [Ros02, p. 52]) to compute
$C_{2}$ $=$ $\frac{q^{2}-1}{q},$
$C_{3}$ $=$ $\frac{(q^{2}+q+1)(q^{2}-1)}{q^{2}}.$
In particular, $C_{3}>2C_{2}$. Hence, in the case of the Hirzebruch surface $X_{2}(1)$ over $K=\mathbb{F}_{q}(T)$, with respect to the height function $H_{L}$ associated with $L=h+f\in\operatorname{Pic}(X_{2}(1))$, our computations show that the component $U_{2}(1)$ in (42) has asymptotically more rational points of large height than the union $\mathbb{P}^{1}\sqcup\mathbb{A}^{1}$ of the other two components.